Write 36x^2-360x-25y^2-100y=100 in standard form. Then state whether the graphs of the equation is a parabola, circle, ellipse, or hyperbola.
Thank You

Answer by fractalier(2101)

From 36x^2 - 360x - 25y^2 - 100y = 100,
factor and complete squares to get
36(x^2 - 10x) - 25(y^2 + 4y) = 100
36(x^2 - 10x + 25) - 25(y^2 + 4y + 4) = 100 + 900 - 100
36(x - 5)^2 - 25(y + 2)^2 = 900
now divide by 900
(x - 5)^2/25 - (y + 2)^2/36 = 1
This is a hyperbola.